# Instance-level Performance Prediction for Long-form Generation Tasks

### v1

@@ 3 Task Formulation
Formalization. Let $\mathcal{X}$ denote the input space (user queries, upstream agent inputs) and $\mathcal{Y}$ the long-form output space (e.g., text, code, etc.) for both LLM generations and gold reference outputs. We assume each task $t\in\mathcal{T}$ has $K_{t}$ continuous, reference-based (RB) evaluation metrics $\mathcal{M}_{t}=\{m_{t,1},\dots,m_{t,K_{t}}\}$, where each metric $m_{t,k}:\mathcal{Y}\times\mathcal{G}\rightarrow\mathbb{R}$ assigns LLM output $y$ a real-valued quality score based on some measure of agreement with reference gold output $g$. At inference, only input $x$ and generation $y$ are observed; gold output $g$ and metric score $m_{t,k}(y,g)$ are not.
@@
Instance-level performance prediction. For each task-specific metric $m_{t,k}$, we seek a predictor
@@
for instance $(x,y)$ that returns: (i) a point estimate $\hat{s}\in\mathbb{R}$ for metric score $m_{t,k}(y,g)$ and (ii) a $(1-\alpha)$ prediction interval $\ell,u]\subseteq\mathbb{R}$ around $\hat{s}$ such that:
$\Pr\!\big{(}m_{t,k}(y,g)\in[\ell,u]\big{)}\geq 1-\alpha,$
@@
Supervision. For each task $t$, assume we have a dataset $\mathcal{D}$ containing $N_{t}$ training instances $\{(x_{i},y_{i},s_{i})\}_{i=1}^{N_{t}}$ for each of $t$’s $K_{t}$ metrics, where $s_{i}=m_{t,k}(y_{i},g_{i})$ denotes score $s_{i}$ for metric $m_{t,k}$, as calculated for generation $y_{i}$ with respect to gold reference $g_{i}$. This setup is both metric-agnostic (allows any metric) and model-agnostic (makes no use of model parameters or decoding internals).
Objective: minimize point-estimation error and also produce calibrated intervals (see Section 5.2).
4 Baseline Methods
Having defined the desired performance predictor $f_{t,k}$ (Section 3), we now describe two baseline methods for instantiating such a predictor.
Confidence-based regression (CE-Reg): Extract unsupervised confidence estimates as features, then learn a regression model to map features to the continuous metric score and a prediction interval. Methods are presented in Sections 4.1 and 4.2.
Reference-Free LLM-as-a-Judge: Prompt an LLM to directly output a point estimate and prediction interval for the target metric, using labeled examples for in-context learning. See Section 4.3.
@@ 4.1 Confidence Estimation (CE)
Verbalized confidence (1S and 2S): Prompting an LLM for its confidence that an output is correct (see Section 2.6). One-step (1S) produces both output and estimated confidence in one invocation, while two-step (2S) first generates the output and then calls the LLM a second time to estimate output confidence. While 1S reduces computation, 2S provides cleaner delineation for benchmarking generation vs. confidence estimation. Note that while 2S permits different LLMs to be used for each step, our experiments use the same LLM for both steps.
@@ 4.2 Confidence-Based Regression (CE-Reg)
Given a dataset $\mathcal{D}=\{(c_{i},s_{i})\}_{i=1}^{N},$ of $N$ training examples, where each $c_{i}$ is the confidence feature (verbalized or consistency-based) for output $y_{i}$ and $s_{i}=m_{t,k}(y_{i},g_{i})$ is the gold score, we fit an auxiliary regression model to learn the conditional distribution $s_{i}\sim p(s\mid c_{i},\theta),$ where $s_{i}=m_{t,k}(y_{i},g_{i})$ is the gold metric score. From the fitted distribution we extract a point estimate $\hat{s}_{i}$ and a $(1-\alpha)$ prediction interval $\ell_{i},u_{i}\,]$ (using $\alpha=0.05$) via
@@
where $F$ is the model’s predictive cumulative distribution function.
@@
CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat. Eccen. EigVal. LexSim. NumSet. 1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.36 0.37 0.38 0.37 0.37 0.38 0.38 0.45 0.43 0.44 Judge(Incon.) 0.28 0.31 0.32 0.31 0.32 0.32 0.32 0.45 0.45 0.44 APIGen CodeBLEU 0.05 0.06 0.06 0.06 0.06 0.06 0.06 0.15 0.14 0.16 LLMEval(A.) 0.16 0.16 0.16 0.15 0.17 0.16 0.16 0.25 0.22 0.23 LLMEval(I.) 0.16 0.16 0.16 0.16 0.16 0.15 0.16 0.24 0.21 0.21 ASQA BERTScore 0.02 0.02 0.02 0.02 0.03 0.02 0.03 0.08 0.08 0.08 BiGGen Judge 0.24 0.30 0.27 0.26 0.31 0.30 0.30 0.31 0.32 0.34 Judge(Func.) 0.26 0.26 0.27 0.25 0.28 0.28 0.28 0.35 0.34 0.35 Judge(Incon.) 0.21 0.21 0.22 0.21 0.21 0.21 0.21 0.27 0.28 0.27 BigCodeBench CodeBLEU 0.08 0.09 0.09 0.08 0.10 0.10 0.10 0.17 0.16 0.17 LLMEval(A.) 0.16 0.17 0.16 0.17 0.17 0.17 0.17 0.22 0.22 0.20 LLMEval(I.) 0.13 0.13 0.14 0.14 0.12 0.13 0.12 0.18 0.17 0.17 Eli5 BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.09 0.09 0.10 Fact-Bio FActScore 0.24 0.33 0.23 0.21 0.33 0.25 0.32 0.33 0.32 0.31 LLMEval(A.) 0.17 0.18 0.17 0.17 0.18 0.18 0.18 0.23 0.22 0.22 LLMEval(I.) 0.15 0.15 0.15 0.15 0.16 0.15 0.15 0.19 0.18 0.18 MED LF QA BERTScore 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.09 0.09 0.09 GEMBA 0.13 0.14 0.13 0.12 0.13 0.14 0.14 0.20 0.20 0.20 Opus-100 BERTScore 0.06 0.07 0.07 0.06 0.08 0.08 0.08 0.12 0.12 0.12 G-Eval(Coh.) 0.13 0.13 0.13 0.13 0.13 0.13 0.13 0.22 0.20 0.19 G-Eval(Con.) 0.11 0.11 0.11 0.11 0.11 0.11 0.11 0.22 0.20 0.19 G-Eval(Flu.) 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.32 0.34 0.36 G-Eval(Rel.) 0.12 0.13 0.13 0.13 0.13 0.13 0.13 0.21 0.20 0.20 Summeval BERTScore 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.06 0.06 0.06 LLMEval(A.) 0.25 0.30 0.26 0.26 0.30 0.30 0.29 0.39 0.40 0.38 LLMEval(I.) 0.22 0.22 0.22 0.22 0.23 0.22 0.23 0.28 0.29 0.27 TruthfulQA BERTScore 0.04 0.04 0.04 0.04 0.04 0.04 0.04 0.15 0.12 0.12 AVERAGE 0.14 0.15 0.15 0.14 0.16 0.15 0.15 0.23 0.22 0.22
Table 1: RMSE Benchmark. We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16- settings. BiGGen-BR = BiGGen-Bench-Results. For RMSE, values lower the better. Judge(Func.) = CodeJudge Functional Correctness. Judge(Incon.) = CodeJudge Inconsistency. LLMEval(A.) = LLM Eval Accuracy. LLMEval(I.) = LLMEval Informativeness.
CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat. Eccen. EigVal. LexSim. NumSet. 1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.36 0.37 0.38 0.37 0.37 0.38 0.38 0.26 0.25 0.26 Judge(Incon.) 0.28 0.31 0.32 0.31 0.32 0.32 0.32 0.33 0.34 0.32 APIGen CodeBLEU 0.05 0.06 0.06 0.06 0.06 0.06 0.06 0.08 0.08 0.08 ASQA LLMEval(A.) 0.16 0.16 0.16 0.15 0.17 0.16 0.16 0.14 0.14 0.14 LLMEval(I.) 0.16 0.16 0.16 0.16 0.16 0.15 0.16 0.14 0.13 0.14 ASQA BERTScore 0.02 0.02 0.02 0.02 0.03 0.02 0.03 0.03 0.03 0.03 BiGGen Judge 0.24 0.30 0.27 0.26 0.31 0.30 0.30 0.20 0.20 0.22 BigCodeBench Judge(Func.) 0.26 0.26 0.27 0.25 0.28 0.28 0.28 0.23 0.23 0.23 Judge(Incon.) 0.21 0.21 0.22 0.21 0.21 0.21 0.21 0.18 0.19 0.18 BigCodeBench CodeBLEU 0.08 0.09 0.09 0.08 0.10 0.10 0.10 0.10 0.10 0.10 LLMEval(A.) 0.16 0.17 0.16 0.17 0.17 0.17 0.17 0.14 0.14 0.13 LLMEval(I.) 0.13 0.13 0.14 0.14 0.12 0.13 0.12 0.11 0.10 0.10 Eli5 BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.02 0.02 0.03 Fact-Bio FActScore 0.24 0.33 0.23 0.21 0.33 0.25 0.32 0.23 0.22 0.20 LLMEval(A.) 0.17 0.18 0.17 0.17 0.18 0.18 0.18 0.14 0.14 0.13 LLMEval(I.) 0.15 0.15 0.15 0.15 0.16 0.15 0.15 0.12 0.11 0.11 MED LF QA BERTScore 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.03 Opus-100 GEMBA 0.13 0.14 0.13 0.12 0.13 0.14 0.14 0.11 0.10 0.10 Opus-100 BERTScore 0.06 0.07 0.07 0.06 0.08 0.08 0.08 0.07 0.07 0.07 G-Eval(Coh.) 0.13 0.13 0.13 0.13 0.13 0.13 0.13 0.14 0.13 0.12 G-Eval(Con.) 0.11 0.11 0.11 0.11 0.11 0.11 0.11 0.13 0.12 0.11 G-Eval(Flu.) 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.27 0.29 0.31 G-Eval(Rel.) 0.12 0.13 0.13 0.13 0.13 0.13 0.13 0.14 0.13 0.12 Summeval BERTScore 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02 LLMEval(A.) 0.25 0.30 0.26 0.26 0.30 0.30 0.29 0.27 0.28 0.26 LLMEval(I.) 0.22 0.22 0.22 0.22 0.23 0.22 0.23 0.19 0.20 0.18 TruthfulQA BERTScore 0.04 0.04 0.04 0.04 0.04 0.04 0.04 0.06 0.05 0.05 AVERAGE 0.07 0.11 0.08 0.07 0.12 0.12 0.12 0.14 0.14 0.14
Table 2: CRPS Benchmark. We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16-shot settings. For CRPS, lower is better.
@@ 4.3 Reference-free LLM-as-a-Judge
Building on prior work (Section 2.6), we define a Reference-free LLM-as-a-Judge (RF-LLMaaJ) baseline for continuous-score performance prediction using in-context learning Vacareanu et al. (2024). It is prompted with: 1) a brief description of the task and the scoring metric; 2) $N$ demonstration examples in form (input, output, gold metric score) of the desired scoring format; 3) instructions to produce a point score and 95% upper/lower interval bound; and 4) A test pair (input, output) for which it must predict a point estimate and prediction interval for the target metric.
5 Benchmark
@@ 5.1 Experimental Setup
Metrics include: BERTScore (Zhang* et al., 2020), CodeBLEU (Ren et al., 2020), and reference-based (RB) LLM-as-a-judge metrics, such as LLM-Eval (accuracy and informativeness) for QA (Xian et al., 2025; Lin and Chen, 2023) and BiGGen (Kim et al., 2025), GEMBA_ref (with references) (Kocmi and Federmann, 2023) for MT, G-Eval (coherence, consistency, fluency, and relevancy) (Liu et al., 2023) for text summarization, FActscore (Min et al., 2023) for fact-checking, and CodeJudge Tong and Zhang (2024) for code generation and function-calling (functional correctness and inconsistency). These LLM judge methods support more robust evaluation by considering diverse aspects of quality (e.g., factual correctness, informativeness, fluency, relevance, etc.).
@@
We also include a “direct assessment” (DA) task (predicting human-annotated quality ratings instead of automatic metric scores, see Section 2.4), for the BiGGen-Bench-Results dataset (Kim et al., 2025). Because this dataset provides human evaluations for its pre-generated LLM outputs, we use both these annotations and outputs for this DA task. However, because the dataset only contain one output per rating, this precludes evaluation of consistency-based methods (see Section 4.1).
@@
LLMs. Llama 3.2 11B (Meta, 2024) is used in both generating outputs and in our baseline predictors (i.e., for consistency and verbalized confidence methods, as well as our RF-LLMaaJ). We set max tokens =769=769, temperature =1=1, and topP =0.9=0.9. To avoid risk of self-enhancement bias (Zheng et al., 2023) in evaluation, we use a different LLM, Gemini-2.0-Flash-Lite, for reference-based (RB) LLM-as-a-judge evaluation metrics.
Our proof-of-concept direct assessment evaluation setting differs by using human quality ratings and existing LLM outputs: Kim et al. (2025)’s GPT-4-0125 (Achiam et al., 2023) generations.
@@
Baselines. Consistency methods are implemented by LM-Polygraph (Fadeeva et al., 2023) and computed over n=3n\!=\!3 outputs, with all $n$ outputs sharing the same consistency score. We implement regression models in scikit-learn and use XGBoost with default hyperparameters. Beta regression is implemented in-house. For regression model selection, we compare RMSE and CRPS results on the dev set. Based on this (Figure 3), we select Random Forest for our test set evaluation.
@@
Regarding scale of supervision, for RF-LLMaaJ in-context learning, we evaluate varying $N\in\{4,8,16\}$ demonstration examples. We only provide up to 16 examples to the RF-LLMaaJ prompt due to high token costs (especially with long-form outputs), and potentially even degrade performance Levy et al. (2024); Modarressi et al. (2025). In contrast, CE-R is trivial to train at scale.
@@
As for relative LLM compute costs of different methods (beyond the output generation call): consistency methods require n1n-1 additional calls (to generate additional outputs for measuring output consistency), verbalized 1S requires no additional calls, verbalized 2S requires one additional call, and RF-LLMaaJ also requires one additional call. In practice, the $n$ outputs generated for consistency would be used to select the output predicted to perform best, so the added compute cost is offset by expected quality improvement in output selection.
Appendix A.2 lists prompts and parsing rules.
5.2 Evaluating Performance Prediction

### v2

@@ 3 PIE Task Formulation
Formalization. Let $\mathcal{X}$ denote the input space (user queries, upstream agent inputs) and $\mathcal{Y}$ the free-form output space (e.g., text, code, etc.) for possible LLM generations, as well as gold reference outputs, if available. We assume each task $t\in\mathcal{T}$ has $K_{t}$ continuous evaluation metrics $\mathcal{M}_{t}=\{m_{t,1},\dots,m_{t,K_{t}}\}$, where each metric $m_{t,k}:\mathcal{Y}\rightarrow\mathbb{R}$ assigns LLM output $y$ a real-valued quality score based on a measure of quality – for example, direct human judgment or agreement with reference gold output $g$. At inference, only input $x$ and generation $y$ are observed; metric score $m_{t,k}(y)$ is not.
@@
Prediction. For each task-specific metric $m_{t,k}$, we seek a predictor defined by:
@@
for instance $(x,y)$ that returns: (i) a point estimate $\hat{s}\in\mathbb{R}$ for metric score $m_{t,k}(y)$ and (ii) a $(1-\alpha)$ prediction interval $\ell,u]\subseteq\mathbb{R}$ around $\hat{s}$ such that:
$\Pr\!\big(m_{t,k}(y)\in[\ell,u]\big)\geq 1-\alpha,$
@@
Supervision. For each task $t$, assume we have a dataset $\mathcal{D}$ containing $N_{t}$ training instances $\{(x_{i},y_{i},s_{i})\}_{i=1}^{N_{t}}$ for each of $t$’s $K_{t}$ metrics, where $s_{i}=m_{t,k}(y_{i})$ denotes the gold score for metric $m_{t,k}$, as assessed for generation $y_{i}$. This setup is both metric-agnostic (allows any metric) and model-agnostic (makes no use of model parameters or decoding internals).
Objective: minimize point-estimation error and produce calibrated intervals (see Section 5.2).
4 Prediction Methods
Having defined the PIE performance predictor $f_{t,k}$ (Section 3), we now describe two such predictors.
Confidence-based Regression (CE-Reg): Extract unsupervised confidence estimates as features, then learn a regression model to map features to the continuous metric score and a prediction interval. Sections 4.1 and 4.2 provide further detail.
Reference-free LLM-as-a-Judge (RF-LLMaaJ): Prompt an LLM to directly output a point estimate and prediction interval for the target metric, using labeled examples for in-context learning. Section 4.3 provides further detail.
@@ 4.1 Confidence Estimation (CE)
Verbalized confidence prompts an LLM for its confidence that an output is correct. One-step (1S) produces both output and estimated confidence in one invocation, while two-step (2S) first generates the output and then calls the LLM a second time to estimate output confidence. While 1S reduces computation, 2S provides cleaner delineation for benchmarking PIE methods on model fixed outputs. See Appendix A.2 for example prompts.
@@ 4.2 Confidence-Based Regression (CE-Reg)
Given a dataset $\mathcal{D}=\{(c_{i},s_{i})\}_{i=1}^{N},$ of $N$ training examples, where each $c_{i}$ is the confidence feature (verbalized or consistency-based) for output $y_{i}$ and $s_{i}=m_{t,k}(y_{i})$ is the gold score, we fit an auxiliary regression model to learn the conditional distribution $s_{i}\sim p(s\mid c_{i},\theta),$ for gold metric score $s_{i}=m_{t,k}(y_{i})$. From the fitted distribution we extract a point estimate $\hat{s}_{i}$ and a $(1-\alpha)$ prediction interval $\ell_{i},u_{i}\,]$ (using $\alpha=0.05$) via
@@
where $F$ is the model’s predictive cumulative distribution function (CDF).
@@ 4.3 Reference-free LLM-as-a-Judge
Building on prior work (Section 2.3), we also compare a reference-free LLM-as-a-Judge (RF-LLMaaJ) for continuous-score estimation using in-context learning Vacareanu et al. (2024). It is prompted with: 1) a brief description of the task and the scoring metric; 2) $N$ demonstration examples in form (input, output, gold metric score); 3) instructions to produce a point score and 95% upper/lower interval bound; and 4) A test pair (input, output) for which it must predict a point estimate and prediction interval for the target metric.
5 Experiments
@@ 5.1 Experimental Setup
Metrics include: BERTScore (Zhang* et al., 2020), CodeBLEU (Ren et al., 2020), and reference-based (RB) LLM-as-a-judge metrics, such as LLM-Eval (accuracy and informativeness) for QA (Xian et al., 2025; Lin and Chen, 2023) and BiGGen (Kim et al., 2025), GEMBA_ref (with references) (Kocmi and Federmann, 2023) for MT, G-Eval (coherence, consistency, fluency, and relevancy) (Liu et al., 2023) for text summarization, FActScore (Min et al., 2023) for fact-checking, and CodeJudge Tong and Zhang (2024) for code generation and function-calling (functional correctness and inconsistency). These LLM judge methods support more robust evaluation by considering diverse aspects of quality, such as factual correctness, informativeness, fluency, relevance, etc.
@@
Direct Assessment (DA) Task. To show PIE’s generality in supporting prediction of both directly observed quantities as well as derived evaluation metrics, we include a proof-of-concept evaluation of predicting observed human quality ratings for the BiGGen-Bench-Results dataset (Kim et al., 2025) (Section 2.2). This dataset provides human ratings for its pre-generated LLM outputs, and we use both its LLM outputs and human ratings. However, because the dataset only contains one output per rating, this precludes evaluation of consistency-based methods (see Section 4.1).
@@
LLMs. Llama 3.2-11B (Meta, 2024) and Gemini-2.0-Flash-Lite (Team et al., 2023) are used to produce output generations. For PIE prediction, whichever LLM was used for generation is then also used for both predictors, CE-Reg and RF-LLMaaJ. We set max tokens =769=769, temperature =1=1, and topP =0.9=0.9. To avoid any self-enhancement bias (Zheng et al., 2023) in ensuring high quality evaluation scores to predict, we use Gemini-2.5-Flash (Team et al., 2023) for reference-based (RB) LLMaaJ evaluation metrics.
The DA task above differs in using existing LLM outputs from prior work: Kim et al. (2025)’s GPT-4-0125 (Achiam et al., 2023) generations.
@@
PIE Predictors. Consistency methods are implemented by LM-Polygraph (Fadeeva et al., 2023) and computed over n=3n\!=\!3 outputs, with all $n$ outputs sharing the same consistency score. We implement regression models in scikit-learn and use XGBoost with default hyperparameters. Beta regression is implemented in-house. For regression model selection, we compare results on the dev set. Based on this (Figure 4), we select Random Forest for our test set evaluation.
@@
For RF-LLMaaJ in-context learning, we vary $N\in\{4,8,16\}$ demonstration examples. We cap at 16 demonstration examples due to high token costs (especially with free-form outputs) and potential to degrade performance Levy et al. (2024); Modarressi et al. (2025). In contrast, CE-Reg is far more efficient to train at scale.
@@
As for relative LLM compute costs of different methods (beyond the output generation call): consistency methods require n1n-1 additional calls (to generate additional outputs for measuring output consistency), verbalized 1S requires no additional calls, verbalized 2S requires one additional call, and RF-LLMaaJ also requires one additional call. In practice, the $n$ outputs generated for consistency would be used to select the output predicted to perform best, so the added compute cost is offset by expected quality improvement in output selection. Appendix A.2 lists prompts and parsing rules.
CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat. Eccen. EigVal. LexSim. NumSet. 1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.22 0.27 0.28 0.25 0.33 0.36 0.34 0.28 0.27 0.26 Judge(Incon.) 0.05 0.09 0.09 0.08 0.11 0.12 0.11 0.14 0.14 0.13 APIGen CodeBLEU 0.02 0.04 0.04 0.03 0.05 0.05 0.05 0.06 0.06 0.06 ASQA LLMEval(A.) 0.14 0.18 0.13 0.13 0.20 0.18 0.19 0.16 0.16 0.16 LLMEval(I.) 0.09 0.13 0.09 0.09 0.14 0.14 0.14 0.13 0.11 0.13 ASQA BERTScore 0.02 0.02 0.02 0.01 0.02 0.02 0.02 0.03 0.02 0.02 BiGGen Judge 0.17 0.26 0.18 0.17 0.26 0.26 0.24 0.25 0.26 0.25 BigCodeBench Judge(Func.) 0.21 0.24 0.22 0.22 0.29 0.29 0.28 0.35 0.37 0.38 Judge(Incon.) 0.03 0.04 0.04 0.04 0.07 0.07 0.06 0.12 0.10 0.09 BigCodeBench CodeBLEU 0.05 0.06 0.05 0.05 0.07 0.07 0.07 0.09 0.08 0.08 LLMEval(A.) 0.15 0.25 0.15 0.15 0.24 0.25 0.23 0.23 0.24 0.23 LLMEval(I.) 0.10 0.14 0.10 0.10 0.13 0.14 0.13 0.14 0.13 0.13 Eli5 BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.02 0.02 0.02 Fact-Bio FActScore 0.10 0.26 0.10 0.11 0.26 0.19 0.26 0.23 0.21 0.22 LLMEval(A.) 0.16 0.25 0.16 0.16 0.25 0.25 0.25 0.23 0.23 0.22 LLMEval(I.) 0.11 0.13 0.11 0.11 0.14 0.14 0.13 0.14 0.13 0.13 MED LF QA BERTScore 0.02 0.02 0.02 0.01 0.02 0.02 0.02 0.03 0.03 0.03 Opus-100 GEMBA 0.10 0.14 0.11 0.11 0.17 0.16 0.16 0.16 0.14 0.14 Opus-100 BERTScore 0.03 0.05 0.04 0.03 0.06 0.06 0.06 0.07 0.07 0.07 G-Eval(Coh.) 0.04 0.05 0.04 0.04 0.05 0.05 0.05 0.08 0.07 0.08 G-Eval(Con.) 0.06 0.08 0.05 0.06 0.08 0.08 0.07 0.10 0.09 0.10 G-Eval(Flu.) 0.01 0.02 0.01 0.01 0.02 0.02 0.02 0.03 0.03 0.03 G-Eval(Rel.) 0.11 0.14 0.11 0.11 0.14 0.14 0.14 0.14 0.14 0.14 Summeval BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.02 0.02 0.02 LLMEval(A.) 0.16 0.29 0.16 0.16 0.31 0.29 0.28 0.30 0.30 0.30 LLMEval(I.) 0.16 0.23 0.17 0.16 0.25 0.24 0.23 0.25 0.25 0.24 TruthfulQA BERTScore 0.02 0.03 0.02 0.02 0.03 0.03 0.03 0.06 0.05 0.05 AVERAGE 0.09 0.13 0.09 0.09 0.14 0.13 0.13 0.14 0.14 0.14
Table 3: CRPS Benchmark. We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16-shot settings. CE-Reg generally scores better (lower) CRPS than RF-LLMaaJ.
5.2 PIE Evaluation